OpenStudy (anonymous):

(7m^2)^3 ??? I'll give medal

OpenStudy (zzr0ck3r):

\((ab)^c=a^cb^c\\(a^b)^c=a^{bc}\)

OpenStudy (anonymous):

is it 7m^5 or 7m^6?

OpenStudy (cp9454):

7m^6

OpenStudy (anonymous):

thank you thats what i thought

OpenStudy (zzr0ck3r):

its none of those

OpenStudy (zzr0ck3r):

using the properties i gave we have \((7m^2)^3=7^3(m^2)^3=343m^6\)

OpenStudy (zzr0ck3r):

Do you understand @Bonnilovescats and @cp9454 ?

OpenStudy (cp9454):

i get it thanks

OpenStudy (anonymous):

yes thanks!